OpenStudy (anonymous):

what is the circumference of the circle x^2+y^2-10y-36=0? a. 38 b. 49 c. 54 d. 125 e. 192

OpenStudy (anonymous):

\[ x^2+y^2-10y-36=0\\ x^2+y^2-10y+25 -25-36=0\\ x^2 +(y-5)^2 = 61=r^2\\ r=\sqrt{61}\\ \]

OpenStudy (anonymous):

\[ 2 \pi r = 2 \pi \sqrt{61}=49.0732 \]

OpenStudy (anonymous):

It seems b is the answer

OpenStudy (anonymous):

Yes it is B.49

OpenStudy (anonymous):

oh, thanks!